OpenStudy (anonymous):

can someone check these answer please. I have an algebra test on this stuff tomorrow, and I just want to know if I'm doing it right! Thanks. MEDAL!! Describe how it was translated from f(x)=x^2

OpenStudy (anonymous):

Give me a second to write the stuff down!

OpenStudy (anonymous):

Questions \[15. f(x)=x^2+3 \] up 3 \[16. f(x)=(x-2)^2\] right 2 \[17.f(x)=x^2-6\] down 6 \[18.f(x)=(x+3)^2\] left 3 \[19. f(x)=x^2-9\] down 9 \[20. f(x)=(x+5)^2\] left 5 \[21. f(x)=x^2+1.5\] left 1.5 \[22.f(x)=(x-2.5)^2\] right 2.5

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

I believe all of them are correct except for number 21, because the 1.5 is not grouped with the x it would go up 1.5 just like you got with the first one. the rest look correct though

OpenStudy (anonymous):

girl you are a life saver thanks I didnt think anyone would respond!! Your amazing!!! A MEDAL FOR YOU!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

happy to help!